OpenStudy (anonymous):

dentify the vertex for the graph of y = -2x^2 - 12x + 1. Answer (-3, 19) (-3, 55) (3, -53) (3, -47)

sam (.sam.):

dentify?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

o identify

sam (.sam.):

\[ax^2-bx+c\] make 'a' 1 -2x^2 - 12x + 1 -2(x^2 +6x)+1 -2(x-3)^2-9(-2)+1 -2(x-3)^3+19

sam (.sam.):

can you find the vertex now?

OpenStudy (anonymous):

can u walk me through it?

sam (.sam.):

ax^2+bx+c=0 The coefficient of x^2 must be 1 everytime you do completing the square, so you must factor 'a' out a(x^2+bx/a)+c=0 then, divide bx/a by 2,then set the x^2 becomes x only,eliminate the x from b/a, then square the whole bracket you ended up with a(x+b/2a)^2+c=0 then, you need to bring the b/2a OUT the brackets and square it, and when you bring it out, it is always negative, be sure that the 'a' will be multiplying the one you bought it out a(x+b/2a)^2-(b/2a)^2(a)+c=0

OpenStudy (anonymous):

k then what?

sam (.sam.):

-2(x-3)^3+19 (x-3)=0 x=3,y=19

OpenStudy (callisto):

I think there're some problems with @.Sam.'s answer.... y = -2x^2 - 12x + 1. = -2(x^2+6x) +1 = -2 (x^2 + 6x +9 - 9) +1 = -2 (x+3)^2 + 18+1 = -2 (x+3)^2 + 19

sam (.sam.):

yes

OpenStudy (anonymous):

isn it -3, 19?

sam (.sam.):

yes

OpenStudy (anonymous):

k thanks

sam (.sam.):

yw